OpenStudy (anonymous):

Solve: x+3squarerootx+2=8

OpenStudy (anonymous):

cube root(x) or 3*sqrt(x)?

OpenStudy (anonymous):

X+3 times square root x+2 = 8(x+2 is under the square root and x+3 is in front of the square root)

OpenStudy (anonymous):

(x+3)sqrt(x+2)=8?

OpenStudy (anonymous):

Approx. x=1.3626

OpenStudy (anonymous):

1.3626426

OpenStudy (anonymous):

I got x=2

OpenStudy (anonymous):

Is it (x+3) times the square root, or 3 times the square root? If it's 3 times, then your answer is correct!

OpenStudy (anonymous):

Yes x+3 times the square root. Thank you!